OpenStudy (anonymous):

I need to use a quadratic formula to solve the equation: 3p(squared)+16p+5=0

OpenStudy (anonymous):

you gotta crunch it out

OpenStudy (anonymous):

a=3, b=16 and c=5, now plug these into the quadratic formula

myininaya (myininaya):

\[3p^2+15p+p+5=0\] \[3p(p+5)+1(p+5)=0\]

myininaya (myininaya):

\[(p+5)(3p+1)=0\]

myininaya (myininaya):

all i did was find 2 factors of a*c that have product a*c and have sum b and then i replaced bp with 15p+1p

OpenStudy (anonymous):

i guess there is no need for the quadratic formula

myininaya (myininaya):

if you want to use you can but i prefer not to use it

myininaya (myininaya):

when possible